OpenStudy (anonymous):

I need to perform the following operation and simplify, but I am stuck: x cubed - 1 / x-1 I am having trouble in factoring the x cubed -1

OpenStudy (anonymous):

x^3-1 = (x-1)(x^2+x+1)

OpenStudy (anonymous):

Can I ask how the factoring is done. The answer makes sense and is correct, but how do you get there? I would like to understand. Can you explain?

OpenStudy (anonymous):

use difference of cubes factor formula A^3 - B^3 = (A-B)(A^2+AB+B^2)

OpenStudy (anonymous):

Thank you. Have been out of algebra for quite a while and am brushing up. Did not remember that formula. Appreciate it!!